A planocovex lens, when silvered in the plane side, behaves like a concave mirror of focal length 30cm. When it is silvered on the convex side,it behaves like a concave mirror of focal length 10 cm. The refractive index of the material of the lens is

Grade:

1 Answers

Arun
6 years ago
Dear Abhinay
1/F = 2/f + 1/f(m)
Plano-convex lens silvered on plane side has f(m) = infinity
1/F = 2/f + 1/ infinity
1/F = 2/f
f = 30*2 = 60 cm
Plano-convex lens silvered on convex side has f(m) = R/2
1/F = 2/f + 2/ R
1/10 = 2/60 + 2/R
R = 30 cm
Now
1/f = (\mu – 1)* R
if we put all value
\mu = 1.5
Regards
Arun (askIITians forum expert)